# Worldline EFT treatment of quadratic and cubic gravity theories

### v1

@@ IV.2 PN-2
$-\frac{(\partial^{i}\partial^{j}\phi\leavevmode\nobreak\ \partial_{i}\partial_{m}\sigma^{k}_{\leavevmode\nobreak\ k}\leavevmode\nobreak\ \partial^{m}\partial_{j}\phi)}{4G\pi}+\frac{3(\partial^{i}\phi\partial^{j}\phi\leavevmode\nobreak\ \partial_{k}\partial_{j}\phi\leavevmode\nobreak\ \partial^{k}\partial_{i}\phi)}{2G\pi}-\frac{(\partial^{j}\partial^{i}\phi\leavevmode\nobreak\ \partial_{i}\partial_{m}\sigma^{k}_{\leavevmode\nobreak\ k}\leavevmode\nobreak\ \partial^{m}\partial_{j}\phi)}{2G\pi}$
$+\frac{3(\partial^{j}\partial^{i}\phi\leavevmode\nobreak\ \partial^{k}\partial_{j}\phi\leavevmode\nobreak\ \partial_{i}\partial_{m}\sigma_{k}^{\leavevmode\nobreak\ m})}{4G\pi}+\frac{3(\partial^{j}\partial^{i}\phi\leavevmode\nobreak\ \partial^{k}\partial_{j}\phi\leavevmode\nobreak\ \partial_{k}\partial_{m}\sigma_{i}^{\leavevmode\nobreak\ m})}{4G\pi}-\frac{3(\partial^{j}\partial^{i}\phi\leavevmode\nobreak\ \partial^{k}\partial_{j}\phi\leavevmode\nobreak\ \partial^{m}\partial_{m}\sigma_{ik})}{4G\pi}$
$+\frac{3(\partial^{i}\partial^{m}\phi\leavevmode\nobreak\ \partial^{k}\partial^{l}\phi\leavevmode\nobreak\ \partial_{i}\partial_{l}\sigma_{km})}{2G\pi}-\frac{3(\partial^{i}\partial^{l}\phi\leavevmode\nobreak\ \partial^{k}\partial^{m}\phi\leavevmode\nobreak\ \partial_{i}\partial_{l}\sigma_{km})}{2G\pi}$ (49)
@@
$-\frac{3(\partial_{j}\partial^{k}\phi(\partial^{m}\partial^{j}A^{i})\partial_{i}\partial_{k}A_{m})}{8G\pi}+\frac{3(\partial^{i}\partial_{l}A_{m}(\partial^{m}\partial^{l}A^{j})\partial_{i}\partial_{j}\phi)}{8G\pi}-\frac{3((\partial_{i}\partial_{l}A^{m})\partial^{i}\partial^{l}A^{j}\partial_{m}\partial_{j}\phi)}{8G\pi}$
@@
We obtain the final result by summing over all the previous contributions at 2 PN and taking an inverse Fourier transform, leading to a change in the two-body effective potential energy due to the Riemann cubed term in the bulk gravitational action compared to GR. At 2 PN, the change is given as follows:
@@
where $r_{12}$ is the inter orbital separation, $v_{i}$ for $i=1,2$ are the velocities of BH 1 and BH 2 respectively, and $a_{i}$ for $i=1,2$ are the accelerations of BH 1 and 2 respectively. Eq. (IV.2) can be rewritten in a form where instead of individual accelerations and velocities, we use relative values. The form is given as follows
@@
$+\frac{108\,G^{2}\,\mu M^{2}}{r_{12}^{8}}\left(\vec{v}_{rel}\cdot\vec{r}_{12}\right)^{2}+\frac{160\,G^{3}\,\mu M^{2}\left(M-2\mu\right)}{7r_{12}^{7}}$
@@ V Conclusions and Discussion
This study explores the two-particle effective potential energy for two modified gravity theories: quadratic gravity and cubic gravity. We integrated the gravitational degrees of freedom using the Worldline Effective Field Theory (WEFT) approach, employing a Kaluza-Klein metric decomposition to simplify the interactions. The analysis showed that higher-order curvature terms—specifically the quadratic Ricci scalar and a particular cubic Riemann tensor contraction—modify the gravitational binding potential at short distances, deviating from the predictions of General Relativity (GR).
For the quadratic gravity case, we observed that the potential energy correction is proportional to the inverse of the fourth power of the separation distance between two black holes. Meanwhile, the cubic gravity term contributed a correction that scaled inversely with the fifth power of the separation. These results suggest that higher-order curvature corrections could significantly alter short-distance gravitational dynamics, potentially offering a means to test these theories in highly compact systems like binary black hole mergers.
@@ VI Acknowledgments
We have partially used the package EFTofPNG [26] in our calculations. We are also indebted to Raj Patil and Jan Steinhoff for important discussions and help regarding the EFTofPNG package.
@@ Appendix A Extra diagrams for cubic gravity at 2 PN
There were eight diagrams that the EFTofPNG package could not solve. We manually calculated them and only tried to gauge the falloff behaviour of some terms. We restricted ourselves to some terms since there was a dramatic proliferation of terms. We found the following
@@
$\begin{split}V_{eff,\alpha}&\propto\int_{k_{1}...k_{5}}d^{3}k_{1}...d^{3}k_{5}\left[\frac{k_{1}^{2}\leavevmode\nobreak\ (k_{2}.k_{3})^{2}\leavevmode\nobreak\ k_{3}^{2}(k_{4}.k_{5})^{2}}{k_{1}^{2}\leavevmode\nobreak\ k_{2}^{2}\leavevmode\nobreak\ k_{3}^{2}\leavevmode\nobreak\ k_{4}^{2}\leavevmode\nobreak\ k_{5}^{2}}\right]e^{ik_{1}.x_{1}}e^{-i(k_{2}+k_{3}+k_{5}).x_{2}}\delta^{3}(k_{1}-k_{2}-k_{3})\delta^{3}(k_{3}-k_{4}-k_{5})\\ &\text{integrate over $k_{1}$ and $k_{3}$}\\ &=\int_{k_{2}k_{4}k_{5}}d^{3}k_{2}\leavevmode\nobreak\ d^{3}k_{4}\leavevmode\nobreak\ d^{3}k_{5}\left[\frac{((k_{2}.(k_{4}+k_{5}))^{2}\leavevmode\nobreak\ (k_{4}.k_{5})^{2}}{k_{2}^{2}\leavevmode\nobreak\ k_{4}^{2}\leavevmode\nobreak\ k_{5}^{2}}\right]e^{i(k_{2}+k_{4}+k_{5}).r}\\ &=\int_{k_{2}k_{4}k_{5}}d^{3}k_{2}\leavevmode\nobreak\ d^{3}k_{4}\leavevmode\nobreak\ d^{3}k_{5}\left[\left(\frac{(k_{2}.k_{4})^{2}(k_{4}.k_{5}))^{2}}{k_{2}^{2}\leavevmode\nobreak\ k_{4}^{2}\leavevmode\nobreak\ k_{5}^{2}}\right)+(k4\Leftrightarrow k_{5})+2\left(\frac{(k_{2}.k_{4})(k_{2}.k_{5})(k_{4}.k_{5})}{k_{2}^{2}\leavevmode\nobreak\ k_{4}^{2}\leavevmode\nobreak\ k_{5}^{2}}\right)\right]\\ &\times e^{i(k_{2}+k_{4}+k_{5}).r}\\ \end{split}$ (53)

### v2

@@ IV.2 PN-2
BIT $=-\frac{(\partial^{i}\partial^{j}\phi\penalty 10000\ \partial_{i}\partial_{m}\sigma^{k}_{\penalty 10000\ k}\penalty 10000\ \partial^{m}\partial_{j}\phi)}{4G\pi}+\frac{3(\partial^{i}\phi\partial^{j}\phi\penalty 10000\ \partial_{k}\partial_{j}\phi\penalty 10000\ \partial^{k}\partial_{i}\phi)}{2G\pi}-\frac{(\partial^{j}\partial^{i}\phi\penalty 10000\ \partial_{i}\partial_{m}\sigma^{k}_{\penalty 10000\ k}\penalty 10000\ \partial^{m}\partial_{j}\phi)}{2G\pi}$
$+\frac{3(\partial^{j}\partial^{i}\phi\penalty 10000\ \partial^{k}\partial_{j}\phi\penalty 10000\ \partial_{i}\partial_{m}\sigma_{k}^{\penalty 10000\ m})}{4G\pi}+\frac{3(\partial^{j}\partial^{i}\phi\penalty 10000\ \partial^{k}\partial_{j}\phi\penalty 10000\ \partial_{k}\partial_{m}\sigma_{i}^{\penalty 10000\ m})}{4G\pi}-\frac{3(\partial^{j}\partial^{i}\phi\penalty 10000\ \partial^{k}\partial_{j}\phi\penalty 10000\ \partial^{m}\partial_{m}\sigma_{ik})}{4G\pi}$
$+\frac{3(\partial^{i}\partial^{m}\phi\penalty 10000\ \partial^{k}\partial^{l}\phi\penalty 10000\ \partial_{i}\partial_{l}\sigma_{km})}{2G\pi}-\frac{3(\partial^{i}\partial^{l}\phi\penalty 10000\ \partial^{k}\partial^{m}\phi\penalty 10000\ \partial_{i}\partial_{l}\sigma_{km})}{2G\pi}$ (67)
The typical vertex factor for the above diagram contains no time derivatives; however, because of the coupling with the worldline, two factors of velocity come into the vertex factor, which, with the propagator and the momentum conservation, is given by the following
$\mathcal{V}=\frac{k_{1}\cdot k_{2}k_{3}^{2}k_{1}\cdot v_{j}k_{2}\cdot v_{k}}{k_{1}^{2}k_{2}^{2}k_{3}^{2}}\delta^{4}(k_{1}+k_{2}+k_{3})\qquad\qquad j,k=1,2$ (68)
Summing over all possible combinations and Fourier integrating the above leads to the following contribution to the effective Lagrangian
$I$ $=\frac{G^{2}m_{1}^{2}m_{2}}{c^{4}}\Bigg[-\frac{12}{r_{12}^{6}}\Bigg(v_{2}^{2}+2v_{1}^{2}\Bigg)-\frac{36}{r_{12}^{8}}\Bigg((v_{2}^{i}x^{i})^{2}+2(v_{1}^{i}x^{i})^{2}\Bigg)\Bigg]$
$+\frac{G^{2}m_{1}m_{2}^{2}}{c^{4}}\Bigg[-\frac{12}{r_{12}^{6}}\Bigg(v_{1}^{2}+2v_{2}^{2}\Bigg)-\frac{36}{r_{12}^{8}}\Bigg((v_{1}^{i}x^{i})^{2}+2(v_{2}^{i}x^{i})^{2}\Bigg)\Bigg]$ (69)
@@
BIT $=-\frac{3(\partial_{j}\partial^{k}\phi(\partial^{m}\partial^{j}A^{i})\partial_{i}\partial_{k}A_{m})}{8G\pi}+\frac{3(\partial^{i}\partial_{l}A_{m}(\partial^{m}\partial^{l}A^{j})\partial_{i}\partial_{j}\phi)}{8G\pi}-\frac{3((\partial_{i}\partial_{l}A^{m})\partial^{i}\partial^{l}A^{j}\partial_{m}\partial_{j}\phi)}{8G\pi}$
@@
A typical vertex factor, propagator, and momentum conservation is given by the following form
$\mathcal{V}=\frac{k_{1}\cdot k_{2}\,k_{1}\cdot k_{3}\,k_{3}\cdot v_{j}\,k_{2}\cdot v_{k}}{k_{1}^{2}k_{2}^{2}k_{3}^{2}}\delta^{4}(k_{1}+k_{2}+k_{3})\qquad\qquad j,k=1,2$ (71)
The above contains two velocity factors, leading to the following contribution to the effective Lagrangian
$\begin{split}I&=\frac{G^{2}}{c^{4}}\Bigg[m_{1}^{2}m_{2}\Bigg(\frac{96(v_{1}^{i}v_{2}^{i})}{r_{12}^{6}}-\frac{144(v_{1}^{i}x^{i})(v_{2}^{i}x^{i})}{r_{12}^{8}}-\frac{12v_{1}^{2}}{r_{12}^{6}}-\frac{36(v_{1}^{i}x^{i})^{2}}{r_{12}^{8}}\Bigg)\\ &+m_{1}m_{2}^{2}\Bigg(\frac{96(v_{1}^{i}v_{2}^{i})}{r_{12}^{6}}-\frac{144(v_{1}^{i}x^{i})(v_{2}^{i}x^{i})}{r_{12}^{8}}-\frac{12v_{2}^{2}}{r_{12}^{6}}-\frac{36(v_{2}^{i}x^{i})^{2}}{r_{12}^{8}}\Bigg)\Bigg]\end{split}$ (72)
@@
Like the $\phi^{3}$ calculations at 1 PN in Sec. IV.1, integrals corresponding to the above diagrams evaluate to zero.
We obtain the final result by summing over all the previous contributions at 2 PN, leading to a change in the two-body effective potential energy due to the Riemann cubed term in the bulk gravitational action compared to GR. At 2 PN, the change is given as follows:
@@
where $r_{12}$ is the inter orbital separation, $v_{i}$ for $i=1,2$ are the velocities of BH 1 and BH 2 respectively, and $a_{i}$ for $i=1,2$ are the accelerations of BH 1 and 2 respectively. Eq. (IV.2) can be rewritten in a form where, instead of individual accelerations and velocities, we use relative values. The form is given as follows
@@
$+\frac{108\,G^{2}\,\mu M^{2}}{r_{12}^{8}}\left(\vec{v}_{rel}\cdot\vec{r}_{12}\right)^{2}+\frac{160\,G^{3}\,\mu M^{2}\left(M-2\mu\right)}{7r_{12}^{7}}+\mathcal{O}\left(\frac{1}{r_{12}^{11}}\right)$
@@
Similar to the quadratic Ricci scalar modified effective potential, the above result is also gauge dependent. The corresponding gauge invariant total binding Energy for circular orbits was found to be the following
$E_{R^{3}}$ $=$ $-\frac{\mu c^{2}x}{2}\left\{1+\left(-\frac{3}{4}-\frac{\nu}{12}\right)x+\left(-\frac{27}{8}+\frac{19}{8}\nu-\frac{\nu^{2}}{24}\right)x^{2}-\beta\left(\frac{1672}{21}+\frac{7744\nu}{21}\right)x^{6}\right\}$
$\beta$ $=$ $\alpha\left(\frac{GM}{c^{2}}\right)^{-4}\qquad\qquad x=\left(\frac{GM\Omega}{c^{3}}\right)^{2/3},$ (76)
Derivation of the above has been given in Appendix C. The above result implies that the actual deviation from GR happens at the sixth PN order, and is consistent with the result of Eq. (2.20) of Ref. [22] since at the leading order $x\propto\frac{1}{r_{12}}$. However, our result has a discrepancy with Eq. (A25) of Ref. [25]. Although both the results are at 6 PN, the coefficient of the 6 PN term differs in our results and theirs. A possible reason could be that Ref. [25] uses 1PN results, while our study uses results up to 2PN.
@@ V Conclusions and Discussion
This study explores the two-particle effective potential energy for two modified gravity theories: quadratic gravity and cubic gravity. We integrated the gravitational degrees of freedom using the Worldline Effective Field Theory (WEFT) approach, employing a Kol-Smolkin metric decomposition to simplify the interactions. The analysis showed that higher-order curvature terms, specifically the particular cubic Riemann tensor contraction, modify the gravitational binding potential at short distances, deviating from the predictions of General Relativity (GR).
We observed that the quadratic gravity theory leads to no change in the effective action compared to GR after integrating out the gravitational degrees of freedom, owing to some field redefinitions that keep the structure of the bulk action the same as in GR. Meanwhile, the cubic gravity term contributed a correction that scaled inversely with the fifth power of the separation. These results suggest that higher-order curvature corrections could significantly alter short-distance gravitational dynamics, potentially providing a means to test these theories in highly compact systems, such as binary black hole mergers. However, the potentials being gauge-dependent cannot be treated as observables. Hence, a gauge-invariant calculation for the total binding energy of a circular orbit was required, which leads to the conclusion that for the cubic modification, the leading-order change in the potential was actually a gauge artefact. For the cubic Riemann case, the actual change happens at the sixth PN order, as was inferred from the gauge invariant binding energy for circular orbits.
@@
The present analysis underscores how the WEFT framework offers a more streamlined and systematic approach to gravitational dynamics than the traditional PN expansion. Whereas PN calculations have evolved over decades to reach high perturbative orders, the WEFT method achieves comparable accuracy through a compact and transparent field-theoretic formulation that integrates out gravitational degrees of freedom using Feynman diagrams. This efficiency facilitates rapid extensions to higher-curvature corrections and modified gravity scenarios while preserving a clear distinction between conservative and radiative sectors. The demonstrated capability of WEFT to handle complex gravitational interactions with speed and clarity suggests its growing relevance for precision studies of relativistic binaries and theoretical explorations beyond General Relativity.
@@ VI Acknowledgments
We have partially used the package EFTofPNG [38] in our calculations. We are indebted to Raj Patil and Jan Steinhoff for important discussions and help regarding the EFTofPNG package. We also want to thank Guillaume Faye for helping us calculate the relativistic centre of mass as a conserved quantity of a PN Lagrangian till 2 PN.
@@ Appendix A Extra diagrams for cubic gravity at 2 PN
There were eight diagrams that the EFTofPNG package could not solve. These eight diagrams, even after letting the calculation for the main file run for several days, did not compute the results nor gave any errors. We believe that the number of terms in calculating the inverse Fourier transform of the amplitudes of the eight diagrams using the results of the Nloop file is computationally expensive for an average machine. We have therefore manually calculated these terms and only tried to gauge the falloff behaviour of some terms. We restricted ourselves to some terms since there was a dramatic proliferation of terms. We found the following
@@
[…]
@@
Appendix B Center of mass for cubic theories
B.1 Center of mass as a conserved quantity
Following Ref. [34], one notices that the effective Lagrangian obtained from a Lorentz invariant gravitational Lagrangian is also itself Lorentz invariant in a PN fashion. Consider the following infinitesimal Lorentz transformation in a general direction with a boost 3-velocity $\beta^{i}$
$\delta_{\beta}x^{i}=\beta^{i}\,t\qquad\delta_{\beta}t=\frac{(x^{i}\beta_{i})}{c^{2}}$ (80)
The above leads to the trajectories $r^{i}_{j}(t)$ for j=1, 2j=1,\,2 transforming as
$\delta_{\beta}r_{j}^{i}$ $=$ $\beta^{i}t-\frac{(\beta^{k}r_{k,j})v_{j}^{i}}{c^{2}}$ (81)
$\delta_{\beta}v_{j}^{i}$ $\equiv$ $\frac{d}{dt}(\delta_{\beta}r_{j}^{i})=\beta^{i}-\frac{(\beta^{k}v_{k,j})v^{i}}{c^{2}}-\frac{(\beta^{k}r_{k,j})a^{i}}{c^{2}}$ (82)
$\delta_{\beta}a_{j}^{i}$ $\equiv$ $\frac{d^{2}}{dt^{2}}(\delta_{\beta}r_{j}^{i})=\mathcal{O}\bigl(\frac{1}{c^{2}}\bigr)$ (83)
For Lagrangians without accelerations, the above transformations corresponds to a changed Lagrangian, which is a total derivative away from the original Lagrangian. The total derivative part or the change in the Lagrangian is given by the following
$\delta_{\beta}L\equiv\frac{d}{dt}\left\{\beta_{i}G^{i}-\sum\limits_{j=1,2}\left[\frac{1}{c^{2}}(p_{j}^{i}v_{i,j})(\beta_{k}r_{j}^{k})\right]\right\}$ (84)
where
$p_{j}^{i}\equiv\frac{\partial L}{\partial v_{i,j}},$ (85)
and $G^{i}$ is the relativistic center of mass vector.
For Lagrangians with accelerations, Eq. (84) changes in the following manner
$\delta_{\beta}L\equiv\frac{d}{dt}\left\{\beta_{i}G^{i}+\sum\limits_{j=1,2}\beta_{i}q_{j}^{i}-\sum\limits_{j=1,2}\left[\frac{1}{c^{2}}(p_{j}^{i}v_{i,j})(\beta_{k}r_{j}^{k})\right]\right\}$ (86)
where
$p^{i}_{j}$ $\equiv$ $\frac{\partial L}{\partial v_{i,j}}-\frac{d}{dt}\left(\frac{\partial L}{\partial a_{i,j}}\right)$ (87)
$q_{j}^{i}$ $\equiv$ $\frac{\partial L}{\partial a_{i,j}}$ (88)
B.2 Modified center of mass for the cubic Riemann theory
The effective Lagrangian for the cubic theory is given by the following
$L=L_{N}+\frac{1}{c^{2}}L_{1PN}+\frac{1}{c^{4}}L_{2PN}$ (89)
where
$L_{N}$ $=$ $\frac{1}{2}m_{1}v_{1}^{2}+\frac{1}{2}m_{2}v_{2}^{2}+\frac{Gm_{1}m_{2}}{r_{12}}$ (90)
$L_{1PN}$ $=$ $\frac{1}{8}m_{1}v_{1}^{4}+\frac{1}{8}m_{2}v_{2}^{4}+\frac{Gm_{1}m_{2}}{r_{12}}\left[\frac{3}{2}\left(v_{1}^{2}+v_{2}^{2}\right)-\frac{7}{2}\left(v_{1}\cdot v_{2}\right)-\frac{1}{2}\left(n_{12}\cdot v_{1}\right)\left(n_{12}\cdot v_{2}\right)\right]$ (91)
$-\frac{G^{2}m_{1}m_{2}\left(m_{1}+m_{2}\right)}{2r_{12}^{2}}$
$L_{2PN}$ $=$ $\frac{1}{16}m_{1}v_{1}^{6}+\frac{1}{16}m_{2}v_{2}^{6}+Gm_{1}m_{2}\left[\frac{7}{4}(a_{2}\cdot v_{1})(n_{12}\cdot v_{1})-\frac{7}{4}(a_{1}\cdot v_{2})(n_{12}\cdot v_{2})\right$. (92)
$+\frac{1}{8}(a_{2}\cdot n_{12})(n_{12}\cdot v_{1})^{2}-\frac{1}{8}(a_{1}\cdot n_{12})(n_{12}\cdot v_{2})^{2}+\frac{7}{8}(a_{1}\cdot n_{12})v_{2}^{2}-\frac{7}{8}(a_{2}\cdot n_{12})v_{1}^{2}$
$\frac{1}{r_{12}}\left\{\frac{7}{8}v_{1}^{2}+\frac{7}{8}v_{2}^{2}+\frac{3}{8}(n_{12}\cdot v_{1})^{2}(n_{12}\cdot v_{2})^{2}-\frac{7}{8}(n_{12}\cdot v_{1})^{2}v_{2}^{2}-\frac{7}{8}(n_{12}\cdot v_{2})^{2}v_{1}^{2}\right$.
$\left.\left.-2v_{1}^{2}(v_{1}\cdot v_{2})-2v_{2}^{2}(v_{1}\cdot v_{2})+\frac{3}{2}(n_{12}\cdot v_{1})(n_{12}\cdot v_{2})(v_{1}\cdot v_{2})+\frac{1}{4}(v_{1}\cdot v_{2})^{2}+\frac{15}{8}v_{1}^{2}v_{2}^{2}\right\}\right]$
$\frac{G^{2}m_{1}m_{2}}{r_{12}^{2}}\left[m_{1}\left\{\frac{7}{2}(n_{12}\cdot v_{1})^{2}+\frac{1}{4}v_{1}^{2}-\frac{7}{2}(n_{12}\cdot v_{1})(n_{12}\cdot v_{2})+\frac{1}{2}(n_{12}\cdot v_{2})^{2}-\frac{7}{4}(v_{1}\cdot v_{2})+\frac{7}{4}v_{2}^{2}\right\}\right$.
$\left.+m_{2}\left\{\frac{1}{2}(n_{12}\cdot v_{1})^{2}+\frac{7}{4}v_{1}^{2}-\frac{7}{2}(n_{12}\cdot v_{1})(n_{12}\cdot v_{2})+\frac{7}{2}(n_{12}\cdot v_{2})^{2}-\frac{7}{4}(v_{1}\cdot v_{2})+\frac{1}{4}v_{2}^{2}\right\}\right]$
[…]
$\left.+\frac{1}{r_{12}}\left[108(n_{12}\cdot v_{1})^{2}-36v_{1}^{2}-216(n_{12}\cdot v_{1})(n_{12}\cdot v_{2})+108(n_{12}\cdot v_{2})^{2}+72(v_{1}\cdot v_{2}-36v_{2}^{2})\right]\right)$
[…]
$\left.\left.\left.\left.+108(n_{12}\cdot v_{2})^{2}+72(v_{1}\cdot v_{2})-36v_{2}^{2}\right]\right)\right\}+\frac{160G^{3}m_{1}m_{2}}{7r_{12}^{7}}\left(m_{1}^{2}+m_{2}^{2}\right)\right]$
Since the modification has both velocity and acceleration, the above will have a non-trivial center of mass modification. Upon applying the transformations (81)-(83) and utilizing Eq. (86), one obtains the following center of mass vector modification (refer to Eq. (B2) in Ref. [34] for the GR relativistic center of mass formula)
$G^{i}_{2PN,R^{3}}=\frac{48\alpha G^{2}m_{1}m_{2}(m_{1}-m_{2})n^{i}_{12}}{r_{12}^{5}}$ (93)
Therefore, the reduced Lagrangian in the center of mass frame is given by
$\mathcal{L}=\mathcal{L}_{N}+\frac{1}{c^{2}}\mathcal{L}_{1PN}+\frac{1}{c^{4}}\mathcal{L}_{2PN}$ (94)
where
$\mathcal{L}_{N}$ $=$ $\frac{v^{2}}{2}+\frac{GM}{r_{12}}$ (95)
$\mathcal{L}_{1PN}$ $=$ $\frac{v^{4}}{8}-\frac{3\,\nu\,v^{4}}{8}+\frac{GM}{r_{12}}\,\left(\frac{\dot{r}^{2}\,\nu}{2}+\frac{3\,v^{2}}{2}+\frac{\nu\,v^{2}}{2}\right)-\frac{G^{2}M^{2}}{2\,r_{12}^{2}}$ (96)
[…]
$\left.\left.+\left(\frac{7}{8}-\frac{5\nu}{4}-\frac{9\nu^{2}}{8}\right)v^{4}\right\}\right]+\frac{G^{2}M^{2}}{r_{12}^{2}}\left[\left(\frac{1}{2}+\frac{41\nu}{8}+\frac{3\nu^{2}}{2}\right)\dot{r}^{2}\right$.
$\left.+\left(\frac{7}{4}-\frac{27\nu}{8}+\frac{\nu^{2}}{2}\right)v^{2}\right]+\frac{G^{3}M^{3}}{r_{12}^{3}}\left(\frac{1}{2}+\frac{15\nu}{4}\right)+\left(\frac{1}{16}-\frac{7\nu}{16}+\frac{13\nu^{2}}{16}\right)v^{6}$
$+\alpha\left[\frac{G^{2}M^{2}}{r_{12}^{5}}\left\{\left(-24-96\nu\right)a_{n}+\frac{1}{r_{12}}\left(108\dot{r}^{2}-36v^{2}\right)\right\}+\frac{G^{3}M^{3}}{r_{12}^{7}}\left(\frac{160}{7}-\frac{320\nu}{7}\right)\right]$
where $a^{i}=a^{i}_{1}-a^{i}_{2}$ is the relative acceleration, $a_{v}=(a\cdot v)$ and $a_{n}=(a\cdot n_{12})$ are the tangential and radial components of the relative acceleration respectively.
Appendix C Gauge invariant binding energy for circular orbits in the cubic Riemann theory
For Lagrangians with accelerations, the Legendre transformation to the Energy is given by
$\mathcal{E}$ $=$ $p_{\phi}\Omega+q_{n}a_{n}$-$\mathcal{L}$ (98)
where $q^{i}$ was defined in Eq. (88) and $q_{n}=(q\cdot n_{12})$. The relative acceleration was found to have the following form in the center of mass frame
$a^{i}$ $=$ $An^{i}+Bv^{i}$ (99)
where $A$ and $B$ are PN functions and are given by
$A_{N}$ $=$ $-\frac{GM}{r_{12}^{2}}$ (100)
$A_{1PN}$ $=$ $\frac{GM}{r_{12}^{2}}\left(\frac{3\dot{r}^{2}\nu}{2}-v^{2}-3\nu v^{2}\right)+\frac{G^{2}M^{2}(4+2\nu)}{r_{12}^{3}}$ (101)
$A_{2PN}$ $=$ $\frac{GM}{r_{12}^{2}}\left(-\frac{15\dot{r}^{4}\nu}{8}+\frac{45\dot{r}^{4}\nu^{2}}{8}+\frac{9\dot{r^{2}}\nu v^{2}}{2}-6\dot{r}^{2}\nu^{2}v^{2}-3\nu v^{4}+4\nu^{2}v^{4}\right)$
$+\frac{G^{2}M^{2}}{r_{12}^{3}}\left(2\dot{r}^{2}+25\dot{r}^{2}\nu+2\dot{r}^{2}\nu^{2}+\frac{13\nu v^{2}}{2}-2\nu^{2}v^{2}\right)-\frac{G^{3}M^{3}}{r_{12}^{4}}\left(9+\frac{87\nu}{4}\right)$
$+\alpha\left[\frac{G^{2}M^{2}}{r_{12}^{7}}\left(-288\dot{r}^{2}-4608\dot{r}^{2}\nu+144v^{2}+576\nu v^{2}\right)-\frac{G^{3}M^{3}}{r_{12}^{8}}\left(256+640\nu\right)\right]$
$B_{1PN}$ $=$ $\frac{GM}{r_{12}^{2}}\left(4\dot{r}-2\dot{r}\nu\right)$ (103)
$B_{2PN}$ $=$ $\frac{GM}{r_{12}^{2}}\left(-\frac{9\dot{r}^{3}\nu}{2}-3\dot{r}^{3}\nu^{2}+\frac{15\dot{r}\nu v^{2}}{2}+2\dot{r}\nu^{2}v^{2}\right)-\frac{G^{2}M^{2}}{r_{12}^{3}}\left(2\dot{r}+\frac{41\dot{r}\nu}{2}-4\dot{r}\nu^{2}\right)$ (104)
$\frac{G^{2}M^{2}\alpha}{r_{12}^{7}}\left(-144\dot{r}+1152\dot{r}\nu\right)$
We obtain the modified Kepler’s third law for the cubic Riemann case by considering $a^{i}=-\Omega^{2}r_{12}n^{i}_{12}$. We obtain the radial separation as a function of $x$ as
$r_{12}=\frac{GM}{c^{2}}\left[\frac{1}{x}-1+\frac{\nu}{3}+\left(\frac{19\nu}{4}+\frac{\nu^{2}}{9}\right)x+\beta\left(\frac{112}{3}+\frac{64\nu}{3}\right)x^{5}\right]$ (105)
where we have scaled $\alpha$ as
$\alpha=\beta\left(\frac{GM}{c^{2}}\right)^{4}$ (106)
Therefore, we after substituting Eq. (105) in Eq. (98), one obtains the following gauge invariant binding Energy for circular orbits for the cubic Riemann modification
$\mathcal{E}=-\frac{c^{2}x}{2}\left\{1-\left(\frac{3}{4}+\frac{\nu}{12}\right)x+\left(-\frac{27}{8}+\frac{19}{8}\nu-\frac{\nu^{2}}{24}\right)x^{2}-\beta\left(\frac{1672}{21}+\frac{7744\nu}{21}\right)x^{6}\right\}$ (107)
which clearly shows that the potential at 2 PN was a gauge artefact and the actual modification comes at 6 PN.